A certain type of bacteria can move at a speed of 220 micrometers per second.
What is the distance,in meters,moved by the bacteria after 12 hours?

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

s = d / t
d = s * t
---
s = 220e-6 meters / second * 3600 seconds / hour = 0.792 meters / hour
Answer:
d = 0.792 * 12 = 9.504 meters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by stanbon(75887)

A certain type of bacteria can move at a speed of 220 micrometers per second.
What is the distance,in meters,moved by the bacteria after 12 hours?
---------------
distance = rate*time = (220(m/10^6))( 1/sec)((12*3600 sec)
-----
= (220*12*3600/10^6) meters
-----
= 9.504 meters
------------------------
Cheers,
Stan H.
========================